A copper wire of 3m and 1mm diameter is subjected to an extension of 5kg. calculate the elongation produced in the wire of Young modulus of elasticity of copper 120GPa.


1
Expert's answer
2021-10-19T15:09:18-0400

The elongation is


"\Delta L=L\frac{\sigma}{E}=L\frac{F}{AE}=\frac{4Lmg}{\pi d^2E},\\\space\\
\Delta L=\frac{4·3·5·(9.8)}{3.14·0.001^2·120·10^9}=1.56·10^{-3}\text{ m}."

About 1.5 mm.